### 31 Mar 2026 v1

@@ Abstract
Next-generation radio astronomy surveys are delivering millions of resolved sources, but scientific exploitation is increasingly limited by robust, scalable morphology analysis across heterogeneous telescopes and imaging pipelines. We present STRADAViT, a self-supervised Vision Transformer (ViT) continued-pretraining framework for learning transferable encoders from radio astronomy imagery. The framework combines mixed-survey data curation, radio astronomy-aware training-view generation, and a ViT-MAE backbone with optional register tokens. It supports reconstruction-only, contrastive-only, and two-stage branches. Our pretraining dataset comprises $512\times 512$ radio astronomy images drawn from four complementary sources (MeerKAT, ASKAP, LOFAR/LoTSS, and SKA). We evaluate transfer with linear probing and fine-tuning on three morphology benchmarks spanning binary and multi-class settings (MiraBest, LoTSS DR2, and Radio Galaxy Zoo). Relative to the ViT-MAE initialization used for continued pretraining, the best two-stage models improve Macro-F1 in all reported linear-probe settings and in two of three fine-tuning settings, with the largest gain on RGZ DR1. Relative to DINOv2, gains are selective rather than universal: the best two-stage models surpass the strongest DINOv2 baseline on LoTSS DR2 and RGZ DR1 under linear probing, and on MiraBest and RGZ DR1 under fine-tuning. A targeted DINOv2-initialized HCL ablation further indicates that the adaptation recipe is not specific to the ViT-MAE starting point. The ViT-MAE-based STRADAViT checkpoint nevertheless remains the preferred release because it combines competitive transfer with substantially lower token count and downstream cost than the DINOv2-based alternative. These results indicate that radio astronomy-aware view generation and staged continued pretraining can provide a stronger domain-adapted starting point than out-of-the-box ViT checkpoints for radio astronomy transfer, especially when representation quality is assessed through linear probing.
@@
{highlights}
Mixed-survey self-supervised dataset spanning MeerKAT, ASKAP, LoTSS, and SKA-derived cutouts
Radio astronomy-aware view generation for sparse and heterogeneous radio astronomy images
Controlled comparison of reconstruction-only, contrastive-only, and two-stage continued pretraining
Two-stage STRADAViT improves on the ViT-MAE starting point in all reported linear-probe settings and in most fine-tuning settings
Domain-adapted transfer is strongest under frozen-backbone evaluation, with selective gains against DINOv2 baselines
@@ 1 Introduction
In mainstream computer vision, transferable representations are increasingly learned via self-supervised pretraining at scale and then adapted to diverse downstream tasks. Vision Transformers (ViTs) build on the Transformer paradigm (Vaswani et al., 2017) and have proven effective when scaled and pretrained appropriately (Dosovitskiy et al., 2021). Self-supervised objectives such as masked image modeling (He et al., 2022) and contrastive or invariance-based learning (Chen et al., 2020; He et al., 2020; Grill et al., 2020; Bardes et al., 2022) enable representation learning without labels, while recent work demonstrates strong ViT features learned purely from self-supervision (Caron et al., 2021; Oquab et al., 2023). A key open question for radio astronomy is how to adapt these pretraining ideas under the constraints of scientific imaging: single-channel data, survey-dependent intensity distributions, and heterogeneous instrumental/imaging systematics.
This paper presents STRADAViT, developed as part of STRADA (Self-supervised Transformers for Radio Astronomy Discovery Algorithms). The framework is organized around three components: (i) a mixed-survey self-supervised pretraining dataset drawn from four complementary radio astronomy image sources, (ii) radio astronomy-aware training-view generation for masked reconstruction and contrastive learning, and (iii) a continued-pretraining framework built on ViT-MAE. The framework supports reconstruction-only, contrastive-only, and two-stage branches.
@@
  • a ViT-MAE-based self-supervised framework with controlled ablations of reconstruction losses, hard-negative-aware contrastive losses, and register tokens, evaluated through reconstruction-only, contrastive-only, and two-stage continued pretraining;
@@
Section 2 reviews the relevant vision and radio-astronomy literature. Sections 3 and 4 describe the pretraining and evaluation datasets. Sections 57 define the model, training branches, and pretraining objectives. Transfer results are then presented under linear probing and fine-tuning. We then present classwise confusion-pattern comparisons against the starting-point baselines, followed by an overall interpretation of the branch comparison and the conclusion.
@@ 2 Literature Survey
This survey covers the transferable-representation literature most directly relevant to the present design choices.
2.1 Backbones for Transferable Visual Representations
@@
Convolutional neural networks remain strong imaging baselines, particularly in residual form (He et al., 2016), but Transformer-based backbones now dominate large-scale representation learning (Vaswani et al., 2017; Dosovitskiy et al., 2021). In specialized imaging domains, backbone choice interacts closely with pretraining objective, dataset composition, and normalization.
2.2 Self-Supervised Learning: Contrastive Invariances and Masked Modeling
@@
Self-supervised learning replaces manual labels with surrogate objectives. Contrastive methods such as SimCLR and MoCo learn invariances across augmented views (Chen et al., 2020; He et al., 2020), whereas related methods modify the handling of negatives or regularize feature statistics (Grill et al., 2020; Bardes et al., 2022). Masked image modeling provides a complementary signal: MAE reconstructs masked content from partial observations and is effective when image statistics differ substantially from those of natural images (He et al., 2022). DINO-style methods and DINOv2 further show that training recipe and architectural details, including register tokens, materially affect transfer (Caron et al., 2021; Oquab et al., 2023). These two objective families motivate the staged design studied here.
2.3 From General-Purpose Pretraining to Domain-Specific Transfer
@@
General-purpose pretraining can yield strong transferable features (Radford et al., 2021), but radio astronomy differs markedly from natural-image domains: images are typically single-channel, have survey-dependent noise and dynamic range, and contain instrument- and pipeline-specific artefacts. Effective transfer therefore depends on representative unlabeled data, domain-appropriate normalization, and augmentations that preserve morphology rather than photographic appearance.
2.4 Radio Astronomy: From Source Finding to Transferable Encoders
@@
Beyond morphology classification, radio astronomy has a long tradition of source finding and characterization under survey-specific artefacts and systematics, as illustrated by CAESAR and related practical pipelines (Riggi et al., 2016, 2019). Recent work extends this toward broad reviews of deep learning for radio classification (Riggi et al., 2024), self-supervised transfer and benchmark studies (Riggi et al., 2024; Cecconello et al., 2025), and multimodal systems (Riggi et al., 2025; Drozdova et al., 2025). Additional SSL studies now cover source classification, FR I/FR II transfer, foundation-style pretraining on Radio Galaxy Zoo, and self-supervised learning on MeerKAT continuum images (Baron Pérez et al., 2025; Buatthaisong et al., 2025; Slijepcevic et al., 2023; Lastufka et al., 2024).
2.5 Radio Morphology Datasets and Benchmark Labels
@@
Radio morphology benchmarks remain anchored in expert or citizen-science labels. LoTSS DR2 provides broad morphology labels and feature flags for bright, extended sources (Horton et al., 2025); MiraBest provides curated FR I/FR II splits (Porter and Scaife, 2023); and Radio Galaxy Zoo operationalizes large-scale citizen-science labeling for more complex multi-class settings (Wong et al., 2024). Across these datasets, labels cover only part of the available data and label spaces are not harmonized across surveys, which makes cross-survey evaluation essential.
@@ 2.6 Positioning Relative to Recent Radio Astronomy SSL Literature
Relative to recent radio astronomy SSL studies (Riggi et al., 2024; Cecconello et al., 2025; Lastufka et al., 2024; Baron Pérez et al., 2025; Buatthaisong et al., 2025; Slijepcevic et al., 2023), STRADAViT emphasizes four points: mixed-survey pretraining, ROI-aware view generation for sparse cutouts, a staged reconstruction-to-contrastive objective, and a controlled ViT-only architectural scope. The aim is cross-survey reuse across multiple downstream label spaces rather than optimization for a single benchmark.
@@ 3 Unsupervised Learning Dataset
Our self-supervised pretraining dataset is constructed to reflect the heterogeneity of contemporary radio astronomy imaging across instruments and imaging pipelines. We assemble $512\times 512$ continuum-image cutouts from four complementary sources: MeerKAT MGCLS DR1, ASKAP continuum images accessed via the CSIRO ASKAP Science Data Archive, LOFAR LoTSS continuum mosaics, and simulated continuum images from the first SKA Science Data Challenge (SDC1) DR1 (Bonaldi et al., 2020).
@@ 3.2 Cutout Normalization
3.3 Dataset Composition
Table 1 summarizes the resulting dataset size after preprocessing and filtering.
@@ 3.3 Dataset Composition
3.4 Online View Generation and Extensibility
All views used for self-supervision are generated on-the-fly during training (Sections 6 and 7), rather than being precomputed. As a consequence, a given cutout is typically encountered under many different ROI selections, crop scales, and augmentations over the course of training, which increases effective data diversity without requiring additional offline curation. This runtime design is also intended to make the pretraining dataset extensible: additional telescope or simulation sources can be incorporated by adding new mosaics (or cutouts) that satisfy the same basic FITS tiling and preprocessing interface, without redesigning survey-specific curation rules.
@@ 4 Evaluation Datasets
We evaluate transfer learning on three public, morphology-focused radio astronomy imaging datasets that span different telescopes, labeling paradigms, and class granularities: MiraBest (Porter and Scaife, 2023), the LoTSS DR2 visual-classification sample of Horton et al. (2025), and Radio Galaxy Zoo DR1 (Wong et al., 2024). Together, these datasets allow us to probe binary and multi-class settings and to test robustness across heterogeneous imaging products.
@@
Figure 3: Example cutouts from the three evaluation datasets after preprocessing with the same pipeline used for pretraining and evaluation (Section 3). Panels show samples of each class for (a) RGZ DR1, (b) MiraBest, and (c) the LoTSS DR2 visual-classification sample of Horton et al. (2025).
@@ 4.1 Common Preprocessing
All evaluation images are treated as single-channel inputs and are standardized using the same per-image ZScale contrast stretch described in Section 3 (mapping to $[0,1]$ and clipping), with NaN/Inf values set to zero. The resulting single-channel image is replicated to three channels and resized to the model input resolution using bicubic interpolation. In downstream evaluation this resolution is model-dependent: ViT-MAE-family models use $224\times 224$ inputs, whereas DINOv2-family models use their native $518\times 518$ inputs. For FITS-based datasets, unreadable files and blank cutouts (e.g., all-zero images) are excluded during indexing.
@@ 4.2 MiraBest (Binary FR I/FR II)
MiraBest provides a curated set of morphologically classified radio galaxies intended for machine learning applications (Porter and Scaife, 2023). The released labels include multiple subtypes and confidence levels. For this work we cast MiraBest as a binary Fanaroff–Riley task by mapping all FR I subtypes to FR I and all FR II subtypes to FR II, while excluding hybrid sources. Unless stated otherwise, we further restrict to confident labels and discard uncertain entries to avoid conflating representation quality with label ambiguity. The images are distributed in a batched format; we apply the same normalization and resizing pipeline as for FITS cutouts to obtain a standardized tensor input.
4.3 LoTSS DR2 Visual-Classification Cutouts (Multi-Class Initial Labels)
Horton et al. (2025) present a large visual-classification effort for bright, resolved LoTSS DR2 sources, including an “initial” class taxonomy and additional morphology and environment flags. We download high-resolution LoTSS DR2 cutouts for each source using the public cutout service, centring on the catalog coordinates. Cutout angular size is computed per source from the reported redshift and physical size when available; otherwise we fall back to a fixed cutout size (5 arcmin) and do not apply additional enlargement (factor =1=1).
To keep the benchmark label space well-defined, we use only the initial class taxonomy (FR I, FR II, Hybrid, Spiral, and Relaxed double) and restrict to sources with exactly one active initial label. Samples with no initial label (or ambiguous multiple initial labels) are excluded. We retain the full set of per-source flags in a JSON sidecar for reproducibility and for possible multi-label extensions.
4.4 Radio Galaxy Zoo DR1 (Six-Class Component/Peak Morphology)
Radio Galaxy Zoo DR1 provides citizen-science classifications for complex radio sources, including a consensus level that can be used to filter for high-confidence labels (Wong et al., 2024). From the FIRST subset, we construct a multi-class benchmark based on the discrete component/peak summary provided in the catalog. Specifically, we derive a class label from the number of radio components and peaks and retain the six most common combinations: 1c1p, 1c2p, 1c3p, 2c2p, 2c3p, and 3c3p. We filter the catalog to consensus level $\mathrm{CL}\geq 0.65$ and download FIRST FITS cutouts via the FIRST cutout service, using an adaptive angular size of 1.5 times the catalog largest angular extent (LAE) and enforcing a minimum cutout side length of 8 pixels. To improve data integrity, cutouts are rejected when the WCS-derived image centre deviates from the target coordinate by more than 1 arcmin.

### 08 Apr 2026 v3

@@ Abstract
Next-generation radio astronomy surveys are delivering millions of resolved sources, but robust and scalable morphology analysis remains difficult across heterogeneous telescopes and imaging pipelines. We present STRADAViT, a self-supervised Vision Transformer (ViT) continued-pretraining framework for learning transferable encoders from radio astronomy imagery. The framework combines mixed-survey data curation, radio astronomy-aware training-view generation, and a ViT-MAE-initialized encoder family with optional register tokens. It supports reconstruction-only, contrastive-only, and two-stage branches. Our pretraining dataset comprises $512\times 512$ radio astronomy cutouts drawn from four complementary sources (MeerKAT, ASKAP, LOFAR/LoTSS, and SKA SDC1 simulated data). We evaluate transfer with linear probing and fine-tuning on three morphology benchmarks spanning binary and multi-class settings (MiraBest, LoTSS DR2, and Radio Galaxy Zoo). Relative to the ViT-MAE initialization used for continued pretraining, the best two-stage models improve Macro-F1 in all reported linear-probe settings and in two of three fine-tuning settings, with the largest gain on RGZ DR1. Relative to DINOv2, gains are selective rather than universal: the best two-stage models achieve higher mean Macro-F1 than the strongest DINOv2 baseline on LoTSS DR2 and RGZ DR1 under linear probing, and on MiraBest and RGZ DR1 under fine-tuning. A targeted DINOv2 initialization ablation further indicates that the adaptation recipe is not specific to the ViT-MAE starting point and that, under the same HCL recipe, the register-based DINOv2 initialization is stronger than the non-register alternative. The ViT-MAE-based STRADAViT checkpoint is retained as the released checkpoint because it combines competitive transfer with substantially lower token count and downstream cost than the DINOv2-based alternative. These results indicate that radio astronomy-aware view generation and staged continued pretraining can provide a stronger domain-adapted starting point than off-the-shelf ViT checkpoints for radio astronomy transfer, especially when representation quality is assessed through linear probing.
@@ 1 Introduction
In mainstream computer vision, transferable representations are increasingly learned via self-supervised pretraining at scale and then adapted to diverse downstream tasks. Vision Transformers (ViTs) build on the Transformer paradigm (Vaswani2017Attention) and have proven effective when scaled and pretrained appropriately (Dosovitskiy2021ViT). Self-supervised objectives such as masked image modeling (He2022MAE) and contrastive or invariance-based learning (Chen2020SimCLR; He2020MoCo; Grill2020BYOL; Bardes2022VICReg) enable representation learning without labels, while recent work demonstrates strong ViT features learned purely from self-supervision (Caron2021DINO; Oquab2023DINOv2). A key open question for radio astronomy is how to adapt these pretraining ideas under the constraints of scientific imaging: single-channel data, survey-dependent intensity distributions, and heterogeneous instrumental/imaging systematics.
This paper presents STRADAViT, a self-supervised transfer framework for radio astronomy Vision Transformers.
@@
  • a controlled continued-pretraining framework initialized from ViT-MAE, with optional register tokens and controlled ablations of reconstruction losses, hard-negative-aware contrastive losses, and reconstruction-only, contrastive-only, and two-stage training branches;
  • a targeted DINOv2 initialization ablation under a fixed contrastive recipe, used to test portability of the adaptation pipeline across materially different ViT starting points;
@@
Section 2 reviews the relevant vision and radio-astronomy literature. Sections 3 and 4 describe the pretraining and evaluation datasets. Sections 57 define the model, training branches, and pretraining objectives, and Section 8 summarizes initialization, optimization, and branch-specific training settings. Section 9 then defines the downstream transfer protocol for linear probing and fine-tuning. Section 10 presents the baseline comparisons, continued-pretraining results, the targeted DINOv2 initialization ablation, and the classwise comparison for the selected release checkpoint, followed by the overall interpretation and the conclusion.
@@ 2 Literature Survey
Convolutional neural networks remain strong imaging baselines, particularly in residual form (He2016ResNet), but Transformer-based backbones now dominate large-scale representation learning (Vaswani2017Attention; Dosovitskiy2021ViT). In specialized imaging domains, backbone choice interacts closely with pretraining objective, dataset composition, and normalization, which motivates keeping the backbone family controlled so that transfer differences can be attributed primarily to the pretraining recipe and view generation.
@@
Self-supervised learning replaces manual labels with surrogate objectives. Contrastive methods such as SimCLR and MoCo learn invariances across augmented views (Chen2020SimCLR; He2020MoCo), whereas related methods modify the handling of negatives or regularize feature statistics (Grill2020BYOL; Bardes2022VICReg). Masked image modeling provides a complementary signal: MAE reconstructs masked content from partial observations and is effective when image statistics differ substantially from those of natural images (He2022MAE). DINO-style methods and DINOv2 further show that training recipe and architectural details, including register tokens, materially affect transfer (Caron2021DINO; Oquab2023DINOv2). These two objective families motivate the staged reconstruction-to-contrastive design studied here.
@@
General-purpose pretraining can yield strong transferable features (Radford2021CLIP), but radio astronomy differs markedly from natural-image domains: images are typically single-channel, have survey-dependent noise and dynamic range, and contain instrument- and pipeline-specific artefacts. Effective transfer therefore depends on representative unlabeled data, domain-appropriate normalization, and augmentations that preserve morphology rather than photographic appearance.
2.1 Radio Astronomy Context and Benchmark Labels
@@
Beyond morphology classification, radio astronomy has a long tradition of source finding and characterization under survey-specific artefacts and systematics, as illustrated by CAESAR and related practical pipelines (Riggi2016CAESAR; Riggi2019CaesarPASA). Recent work extends this toward broad reviews of deep learning for radio classification (Riggi2024DeepLearningRadio), self-supervised transfer and benchmark studies (Riggi2024PASAContrastive; Cecconello2025RadioSSLBenchmark), and multimodal systems (Riggi2025RadioLLaVA; Drozdova2025VLMRadio). Additional SSL studies now cover source classification, FR I/FR II transfer, foundation-style pretraining on Radio Galaxy Zoo, and self-supervised learning on MeerKAT continuum images (BaronPerez2025AandASSL; Buatthaisong2025MNRASSelfSupervisedFR; Slijepcevic2024RGZFoundation; Lastufka2024MeerKATSSL).
@@
Radio morphology benchmarks remain anchored in expert or citizen-science labels. LoTSS DR2 provides broad morphology labels and feature flags for bright, extended sources (Horton2025LoTSSDR2Morphology); MiraBest provides curated FR I/FR II splits (PorterScaife2023MiraBest); and Radio Galaxy Zoo operationalizes large-scale citizen-science labeling for more complex multi-class settings (RGZDR1Zenodo14195049). Across these datasets, labels cover only part of the available data and label spaces are not harmonized across surveys, which makes cross-survey evaluation essential.
@@ 2.2 Positioning Relative to Recent Radio Astronomy SSL Literature
Relative to recent radio astronomy SSL studies (Riggi2024PASAContrastive; Cecconello2025RadioSSLBenchmark; Lastufka2024MeerKATSSL; BaronPerez2025AandASSL; Buatthaisong2025MNRASSelfSupervisedFR; Slijepcevic2024RGZFoundation), STRADAViT emphasizes four points: mixed-survey pretraining, ROI-aware view generation for sparse cutouts, a staged reconstruction-to-contrastive objective, and a controlled ViT-only architectural scope. The aim is cross-survey reuse across multiple downstream label spaces rather than optimization for a single benchmark.
@@ 3 Unsupervised Learning Dataset
Our self-supervised pretraining dataset is constructed to reflect the heterogeneity of contemporary radio astronomy imaging across instruments and imaging pipelines. We assemble $512\times 512$ continuum-image cutouts from four complementary sources: MeerKAT MGCLS DR1, ASKAP continuum images accessed via the CSIRO ASKAP Science Data Archive, LOFAR LoTSS continuum mosaics, and simulated continuum images from the first SKA Science Data Challenge (SDC1) DR1 (Bonaldi2020). These sources cover different angular resolutions, noise properties, imaging artefacts, and source morphologies, so the pretraining dataset is intended to expose the model to heterogeneous radio structure rather than to a single survey regime. Table 1 summarizes the resulting dataset size after preprocessing and filtering.
@@ 3.2 Cutout Normalization
3.3 Online View Generation
All views used for self-supervision are generated on-the-fly during training (Sections 6 and 7), rather than being precomputed. A given cutout is therefore typically encountered under many different ROI selections, crop scales, and augmentations over the course of training, which increases effective data diversity without additional offline curation. The same runtime design also makes the pretraining dataset extensible, since additional telescope or simulation sources can be incorporated by adding new mosaics (or cutouts) that satisfy the same FITS tiling and preprocessing interface.
@@ 4 Evaluation Datasets
We evaluate transfer learning on three public radio astronomy imaging datasets that span different telescopes, labeling paradigms, and class granularities: MiraBest (PorterScaife2023MiraBest), the LoTSS DR2 visual-classification sample of Horton2025LoTSSDR2Morphology, and Radio Galaxy Zoo DR1 (RGZDR1Zenodo14195049). Together, these datasets probe compact binary morphology, richer multi-class morphology, and component/peak complexity under heterogeneous imaging products.
@@
Figure 3: Example cutouts from the three evaluation datasets after preprocessing with the same pipeline used for pretraining and evaluation (Section 3). Panels show samples of each class for (a) RGZ DR1, (b) MiraBest, and (c) the LoTSS DR2 visual-classification sample of Horton2025LoTSSDR2Morphology.
@@ 4.1 Common Preprocessing
All evaluation images are treated as single-channel inputs and are standardized using the same per-image ZScale contrast stretch described in Section 3 (mapping to $[0,1]$ and clipping), with NaN/Inf values set to zero. The resulting single-channel image is replicated to three channels and resized to the model input resolution using bicubic interpolation. In downstream evaluation this resolution is model-dependent: ViT-MAE-family models use $224\times 224$ inputs, whereas DINOv2-family models use their native $518\times 518$ inputs.
@@ 4.2 MiraBest (Binary FR I/FR II)
MiraBest provides a curated set of morphologically classified radio galaxies intended for machine learning applications (PorterScaife2023MiraBest). The released labels include multiple subtypes and confidence levels. For this work we cast MiraBest as a binary Fanaroff–Riley task by mapping all FR I subtypes to FR I and all FR II subtypes to FR II, while excluding hybrid sources. Unless stated otherwise, we further restrict to confident labels and discard uncertain entries to avoid conflating representation quality with label ambiguity. We apply the same normalization and resizing pipeline as for FITS cutouts to obtain a standardized tensor input.
4.3 LoTSS DR2 (Multi-Class Initial Labels)
Horton2025LoTSSDR2Morphology present a large visual-classification effort for bright, resolved LoTSS DR2 sources, including an “initial” class taxonomy and additional morphology and environment flags. We download high-resolution LoTSS DR2 cutouts for each source using the public cutout service, centring on the catalog coordinates. Cutout angular size is computed per source from the reported redshift and physical size when available; otherwise we fall back to a fixed cutout size (5 arcmin) and do not apply additional enlargement (factor =1=1).
To keep the benchmark label space well-defined, we use only the initial class taxonomy (FR I, FR II, Hybrid, Spiral, and Relaxed double) and restrict to sources with exactly one active initial label. Samples with no initial label (or ambiguous multiple initial labels) are excluded.
4.4 Radio Galaxy Zoo DR1 (Multi-Class Component/Peak Morphology)
Radio Galaxy Zoo DR1 provides citizen-science classifications for complex radio sources, including a consensus level that can be used to filter for high-confidence labels (RGZDR1Zenodo14195049). From the FIRST subset, we construct a multi-class benchmark based on the discrete component/peak summary provided in the catalog. Specifically, we derive a class label from the number of radio components and peaks and retain the six most common combinations: 1c1p, 1c2p, 1c3p, 2c2p, 2c3p, and 3c3p. We filter the catalog to consensus level $\geq 0.65$ and download cutouts via the FIRST cutout service, using an adaptive angular size of 1.5 times the catalog largest angular extent (LAE) and enforcing a minimum cutout side length of 8 pixels. To improve data integrity, cutouts are rejected when the WCS-derived image centre deviates from the target coordinate by more than 1 arcmin.